# A novel characterization of structures in smooth regression curves: from a viewpoint of persistent homology

### v3

@@ 3 Problem formulation and main results
Note that $\{D_{t}\}_{t\in\mathbb{R}}$ is a nested family of super-level sets as $t$ decreases from $\infty$ to $\infty$- ∞ that is for any $t_{1}<t_{2}$, we have $D_{t_{2}}\subset D_{t_{1}}$. We use persistent homology for the filtration $\{D_{t}\}_{t\in\mathbb{R}}$ to summarize the geometric features of $m_{1}$. Also note that $m_{1}$ is unknown since $m$ is unknown, so we estimate persistent homology of $m_{1}$ from the observed data $\mathcal{X}_{n}\triangleq\{(X_{1},Y_{1}),\ldots(X_{n},Y_{n})\}$. The common procedure is to use an estimate $\hat{f}$ to estimate the persistent homology of an unknown function $f$.
@@
    Note that for any $\epsilon\in(0,|t|)$, we have $\hat{D}_{t+\epsilon}(n,h)\subset\hat{D}_{t-\epsilon}(n,h)$. This gives an inclusion map from $\hat{D}_{t+\epsilon}(n,h)\text{ to }\hat{D}_{t-\epsilon}(n,h)$.
@@
    Now, estimate the homology using $\hat{H}_{k}(t_{i},\epsilon;n)$ for every $i\in\mathbb{Z}$, and compute the persistent homology for the following discrete filtration $\hat{\mathcal{D}}^{\epsilon}\triangleq\{\hat{D}_{t_{i}}(n,h)\}_{i\in\mathbb{Z}}$. We denote this estimator of persistent homology as $\widehat{PH}^{\epsilon}_{k}(f)$.
@@
where $K$ is a kernel function, $h\equiv h(n)$ is kernel bandwidth and $K_{h}(x)=K(x/h)$. We require point-wise consistency of $\hat{m}_{n}(x)$ for theoretical considerations in this paper. Therefore, we assume the following regularity conditions on the Kernel function:
@@
(A.2) The support of kernel $K$ is a set $S\subseteq[-1,1]$ such that $K(x)=K(-x)$ for all $x\in S$.
@@
Then under the Assumptions (A.1) - (A.4), Theorem 2.1 of [31] states that if $x\in\mathbb{R}$ with $p(x)\neq 0$ is a continuous point of $p$ (marginal density of X), $m$ (the regression function), and $f$ (joint density of (X,Y)) for each fixed $y\in\mathbb{R}$, then
$\hat{m}_{n}(x)\xrightarrow{a.s}m(x)\text{ as }h\xrightarrow{}0\text{ and }n\xrightarrow{}\infty\text{ such that }nh\xrightarrow{}\infty.$ (3.8)
@@
where ${K_{1,h}(x-X_{i})=\frac{\mathrm{d}}{\mathrm{d}x}K_{h}(x-X_{i})}$ assuming that first order derivative of the kernel function $K$ exists. In addition, we assume the following conditions on the first order derivative of K denoted as $K_{1}$:
(A.5) For any real number $\tau>0$ we have $\left|K_{1}(x)\right|\leq\tau$ for all $x\in S$ and $K_{1}(0)=0$.
@@
(B.1) The marginal density of X denoted as p is tame (see, Definition 2.3) and bounded.
(B.2) The support of p is a compact set $\mathcal{X}$, and $p_{min}=\inf_{x\in\mathcal{X}}p(x)>0$, and $p_{max}=\sup_{x\in\mathcal{X}}p(x)$.
(B.3) The response variable satisfies $|Y|\leq Y_{max}$ almost surely, $Y_{max}$ is a positive real number.
(B.4) The regression function $m$ is continuous, and $|m(x)|\leq M$ for all $x\in\mathbb{R}$, where $M>0$.
(B.5) The derivative of the regression function $m_{1}$ is tame and continuous on a compact domain, and $|m_{1}(x)|\leq M_{1}$ for all $x\in\mathbb{R}$, where $M_{1}$ is a positive constant.
@@
If $h\equiv h(n)\xrightarrow{}0$ as $n\xrightarrow{}\infty$ such that $nh^{2}\xrightarrow{}\infty$, then under the Assumptions (A.1) - (A.5) and (B.1) -( B.5), we have
$\mathds{P}\left(\delta_{B}\left(\widehat{PH}_{k}^{\epsilon}\left(m_{1}\right),PH_{k}\left(m_{1}\right)\right)\leq 5\epsilon\right)\geq 1-3N_{\epsilon}ne^{-{C_{\epsilon/2}nh^{2}}},$
where k is any arbitrary degree of homology, $\delta_{B}$ is the bottleneck distance (see, Definition 2.5), $N_{\epsilon}$ is defined in Equation 3.6, and
$C_{\epsilon}=\displaystyle{\frac{2\epsilon^{2}\delta^{2}p_{min}^{2}}{\left[\tau\left(Y_{max}+M\right)+(M_{1}+\epsilon)\right]^{2}}}.$
[…]
@@
Note that, the constant $C_{\epsilon}$ must be positive in order to have a valid lower bound of the probability in Theorem 3.1. Since if $C_{\epsilon}=0$, then the sequence $01-3N_{\epsilon}ne^{-{C_{\epsilon/2}nh^{2}}}=1-3N_{\epsilon}n<$ 0. Furthermore, as $n\xrightarrow{}\infty$ the sequence goes to $\infty$ implying that the proposed procedure is not consistent. For this reason, we need to assume that $\delta>0$ as precisely defined in Assumption (A.4).
Theorem 3.1 states that the estimator $\widehat{PH}_{k}^{\epsilon}\left(m_{1}\right)$ is consistent up to the precision of $5\epsilon$. However, one can choose $\epsilon$ arbitrarily small to acquire higher precision. Note that $\epsilon\in(0,|t|/2)$ such that the condition in Definition 3.1 is satisfied. We have assumed that the levels are $\epsilon$ regular (see, Definition 3.1). Therefore, such an $\epsilon$ always exists, and we need not choose the precise value of $\epsilon$, instead choosing any lower bound will suffice. Note that the smaller the value of $\epsilon$, the smaller the value of $C_{\epsilon}$ will be, and the convergence of $\widehat{PH}_{k}^{\epsilon}\left(m_{1}\right)$ to $PH_{k}\left(m_{1}\right)$ will be slower.
[…]
@@ 4 Statistical significance of local structures
In the following sections, we apply Theorem 3.1 in identifying the local structures such as monotonicity, convexity, and modality in smooth regression curves. One major challenge in inferring local structures of regression functions from a finite data set is assessing their statistical significance. A quantitative statistical assessment is important to determine whether the observed local structures are signals or artifacts of sampling noise. In the present context, local structures are derived from the 0-dimensional estimated persistence barcodes computed using the procedure given in Section 3 whose validity is guaranteed by Theorem 3.1. Therefore, we translate the statistical significance of the observed local structures in terms of the statistical significance of the 0-dimensional topological features (connected components) in the estimated barcode.
@@
We propose a measure of statistical significance of the observed 0-dimensional topological features in the estimated barcodes calculated from the procedure described in Section 3. We translate the statistical significance of the 0-dimensional topological features in the estimated barcodes in terms of the statistical significance of the critical values of the estimated derivative. Since for the super level-set filtrations of the Morse and tame functions, topological features in the 0-dimensional persistence barcodes are born and die at the critical values of the functions. Also, note that, for any $k\geq 0$, the birth and the death time of a $k$-dimensional topological feature completely characterize the topological feature in the $k$-dimensional barcode. Consequently, it is equivalent to having the statistical significance of the birth and the death times of topological features in the barcode to have the statistical significance of the topological features in the estimated barcode. Hence, we measure the statistical significance of the critical values of the estimated derivative to measure the statistical significance of the 0-dimensional topological features in the estimated barcode. To do this, let us simplify the expression of the estimated derivative given in Equation (3.9), We have
$\hat{m}_{1,n}(x)=\frac{\sum\limits_{i=1}^{n}K_{1,h}\left(x-X_{i}\right)\left(Y_{i}-\hat{m}_{n}(x)\right)}{\sum\limits_{i=1}^{n}K_{h}\left(x-X_{i}\right)}=\sum\limits_{i=1}^{n}\omega_{i}\left(x\right)r_{i}(x),$ (4.1)
where ${\omega_{i}(x)=\frac{K_{1,h}\left(x-X_{i}\right)}{\sum\limits_{i=1}^{n}K_{h}\left(x-X_{i}\right)}}$ are weights to the $i$-th residual $r_{i}(x)=Y_{i}-\hat{m}_{n}(x)$, $i=1,\ldots,n$.
Note that, from Equation (4.1), for any fix $x_{0}\in\mathbb{R}$, the value $\hat{m}_{1,n}(x_{0})$ is the weighted sum of the residuals $r_{i}(x_{0}),i=1,\ldots,n$. Given this fact, we declare $\hat{m}_{1,n}(x_{0})$ to be significantly different from noise (statistically significant) if the residuals $r_{i}(x_{0})$ that have larger values contribute less in Equation 4.1 compared to those that have smaller values. In other words, we declare the value $\hat{m}_{1,n}(x_{0})$ to be statistically significant if the larger weights are assigned in the corresponding Equation 4.1 to the residuals that have smaller values. In particular, if we observe a significantly decreasing trend between the residuals $r_{i}(x_{0})$, and the weights $\omega_{i}(x_{0})$ for all $i=1,\ldots,n$, we say that the value $\hat{m}_{1,n}(x_{0})$ is statistically significant. This phenomenon can be captured by fitting a linear regression model with $r_{i}(x_{0})$ as the response variable and $\omega_{i}(x_{0})$ as a regressor, for $i=1,\ldots,n$. Consider the following linear regression model:
$r_{i}(x_{0})=\beta_{0}+\beta_{1}\omega_{i}(x_{0})+\eta_{i},i=1,\ldots,n$ (4.2)
where $\beta_{0}$ and $\beta_{1}$ are unknown parameters and $\eta_{i}$ is unobserved random error in the model with the usual error assumptions. We say that the value $\hat{m}_{1,n}(x_{0})$ is statistically significant if the estimate of $\beta_{1}$ is negative and statistically significant.
Now, we illustrate the procedure described above. Suppose that we observe a barcode with $p$ 0-dimensional topological features, say $\mathcal{B}_{0}=\{[a_{1},b_{1}],\ldots,[a_{p},b_{p}]\}$, $p$ is a finite positive integer, from the super-level set filtration of the estimated derivative $\hat{m}_{1,n}$ using the procedure described in Section 3. Then for each 0-dimensional feature $[a_{i},b_{i}]$ in $\mathcal{B}_{0}$, there exists critical points, $x_{i}^{a},x_{i}^{b}\in\mathbb{R}$, of $\hat{m}_{1,n}$ such that $\hat{m}_{1,n}(x_{i}^{a})=a_{i}$, and $\hat{m}_{1,n}(x_{i}^{b})=b_{i}$, respectively, for $i=1,\ldots,p$. We say that a feature $[a_{i},b_{i}]$ in $\mathcal{B}_{0}$ is statistically significant if both the values $\hat{m}_{1,n}(x_{i}^{a})=a_{i}$, and $\hat{m}_{1,n}(x_{i}^{b})=b_{i}$ are statistically significant; otherwise, we say that the feature is statistically insignificant, for $i=1,\ldots,p$. We demonstrate this procedure on simulated and real data in the following sections.

### v4

@@ 3 Problem formulation and main results
Note that $\{D_{t}\}_{t\in\mathbb{R}}$ is a nested family of super-level sets as $t$ decreases from $\infty$ to -$\infty$ that is for any $t_{1}<t_{2}$, we have $D_{t_{2}}\subset D_{t_{1}}$. We use persistent homology for the filtration $\{D_{t}\}_{t\in\mathbb{R}}$ to summarize the geometric features of $m_{1}$. Also note that $m_{1}$ is unknown since $m$ is unknown, so we estimate persistent homology of $m_{1}$ from the observed data $\mathcal{X}_{n}\triangleq\{(X_{1},Y_{1}),\ldots(X_{n},Y_{n})\}$. The common procedure is to use an estimate $\hat{f}$ to estimate the persistent homology of an unknown function $f$.
@@
[…]
@@
    Now, estimate the homology using $\hat{H}_{k}(t_{i},\epsilon;n)$ for every $i\in\mathbb{Z}$, and compute the persistent homology for the following discrete filtration $\hat{\mathcal{D}}^{\epsilon}\triangleq\{\hat{D}_{t_{i}}(n,h)\}_{i\in\mathbb{Z}}$. This estimator of persistent homology is denoted as $\widehat{PH}^{\epsilon}_{k}(f)$.
@@
where $K$ is a kernel function, $h\equiv h(n)$ is kernel bandwidth and $K_{h}(x)=K(x/h)$. We require the following regularity conditions on the Kernel function:
@@
(A.2) The support of kernel $K$ is a set $S\subseteq[-1,1]$ such that K(x)=K(x)K(x)=K(-x) for all $x\in S$.
@@
where ${K_{1,h}(x-X_{i})=\frac{\mathrm{d}}{\mathrm{d}x}K_{h}(x-X_{i})}$ assuming that the first order derivative of the kernel function $K$ exists. In addition, we restrict the class of kernel functions satisfying Assumptions (A.1)-(A.4) to the following class of kernel functions:
(A.5) For any real number $b>0$, the class of kernel functions is defined as:
$\left\{K:\frac{\mathrm{d}}{\mathrm{d}x}K\left(\frac{x}{b}\right)=\frac{x}{b^{2}}\psi\left(\frac{x}{b}\right)K\left(\frac{x}{b}\right)\right\},$
where $\psi$ is a function such that $\tau_{1}\leq\psi(x)\leq\tau_{2}$, for all $x\in\mathbb{R}$. Here, $\tau_{1},\tau_{2}\in\mathbb{R}$ such that $\tau_{1}\leq\tau_{2}$.
Remark 3.1
The standard Gaussian and Cauchy kernels satisfy Assumption (A.5) with $\psi(x)=-1$, and $\psi(x)=2\left(1+x^{2}\right)^{-1}$, respectively, for $x\in\mathbb{R}$.
@@
(B.1) The marginal density of $X$ denoted as p is tame (see, Definition 2.6) and bounded. Moreover, $X_{min}<X<X_{max}$ almost surely, where $X_{min}$ and $X_{max}$ are real numbers.
(B.2) The support of p is a compact set $\mathcal{X}$ such that, $p_{min}=\inf_{x\in\mathcal{X}}p(x)>0$, and $p_{max}=\sup_{x\in\mathcal{X}}p(x)$.
(B.3) The response variable satisfies $Y_{min}<Y<Y_{max}$ almost surely, where $Y_{min}$ and $Y_{max}$ are real numbers.
(B.4) The regression function $m$ is continuous, and $\gamma<m(x)<M$ for all $x\in\mathbb{R}$, where $\gamma$ and $M$ are real numbers.
(B.5) The derivative of the regression function $m_{1}$ is tame and continuous on a compact domain, and $\gamma_{1}<m_{1}(x)<M_{1}$ for all $x\in\mathbb{R}$, where $\gamma_{1},M_{1}\in\mathbb{R}$.
@@
If $h\equiv h(n)\xrightarrow{}0$ as $n\xrightarrow{}\infty$ such that $nh^{6}\xrightarrow{}\infty$, then under the Assumptions A.1 - A.5 and B.1 - B.5, we have
$\mathds{P}\left(\delta_{B}\left(\widehat{PH}_{k}^{\epsilon}\left(m_{1}\right),PH_{k}\left(m_{1}\right)\right)\leq 5\epsilon\right)\geq 1-3M_{\epsilon}ne^{-{C_{\epsilon/2}nh^{6}}},$ (3.9)
where k is any arbitrary degree of homology, $\delta_{B}$ is the bottleneck distance (see, Definition 2.8), $M_{\epsilon}$ is a positive integer defined as $M_{\epsilon}=\min\{i:t_{i}\geq\gamma_{1}\}$, $t_{i}$ is defined in Equation (3.6), and $C_{\epsilon}$ is defined as
$C_{\epsilon}=\displaystyle{\frac{8\delta^{2}p_{min}^{2}(\gamma_{1}+\epsilon)^{2}}{(X_{max}-X_{min})^{2}\left[\tau_{2}(Y_{max}-Y_{min}\delta)+\tau_{1}Y_{min}\delta(1-\delta)\right]^{2}}}\text{, provided }\gamma_{1}\neq-\epsilon.$ (3.10)
Here all the involved constants are precisely defined in Assumptions (A.1)-(A.5) and (B.1)-(B.5).
@@
Note that the constant $C_{\epsilon}$ must be positive in order to have a valid lower bound of the probability in Theorem 3.1. Since if $C_{\epsilon}=0$, then the sequence $1-3M_{\epsilon}ne^{-{C_{\epsilon/2}nh^{6}}}=1-3M_{\epsilon}n<0$. Furthermore, as $n\xrightarrow{}\infty$, the sequence goes to -$\infty$, implying that the proposed procedure is not consistent. For this reason, we need to assume that $\delta>0$ as precisely defined in Assumption (A.4).
@@ 4 Statistical significance of local structures
In the following sections, we apply Theorem 3.1 in identifying the local structures such as monotonicity, convexity, and modality in smooth regression curves. One major challenge in inferring local structures of regression functions from a finite data set is assessing their statistical significance. A quantitative statistical assessment is required to determine whether the observed local structures are signals or artifacts of sampling noise. In the present context, local structures are derived from the 0-dimensional estimated persistence barcodes computed using the procedure given in Section 3 whose validity is guaranteed by Theorem 3.1. Therefore, we translate the statistical significance of the observed local structures in terms of the statistical significance of the 0-dimensional topological features (connected components) in the estimated barcode.
@@
We propose a measure of statistical significance of the observed 0-dimensional topological features in the estimated barcodes calculated from the procedure described in Section 3. We provide a significance measure for the topological features in the observed barcode using Theorem 3.1. For $\alpha\in(0,1)$, a $(1-\alpha)$-confidence set using Theorem 3.1 can be constructed as:
$\mathcal{C}_{n}:=\left\{\mathcal{B}:\delta_{B}\left(\mathcal{B},\hat{\mathcal{B}}\right)\leq 5z_{n,\alpha}\right\},$ (4.1)
where $\mathcal{B}$ is an arbitrary barcode, $\hat{\mathcal{B}}$ is a barcode estimated using the procedure given in Section 3, and $z_{n,\alpha}$ solves the following:
$3M_{z_{n,\alpha}}n\exp\left({-nh^{6}C_{z_{n,\alpha}/2}}\right)=\alpha.$ (4.2)
Here $M_{z_{n,\alpha}}$ and $C_{z_{n,\alpha}}$ are defined by plugging $z_{n,\alpha}$ in place of $\epsilon$ in $M_{\epsilon}$ and $C_{\epsilon}$ (see Equation (3.10)), respectively.
Note that $\mathcal{C}_{n}$ contains the true barcode $\mathcal{B}_{0}$ with probability $1-\alpha$. That is
$\mathbb{P}\left(\mathcal{B}_{0}\in\mathcal{C}_{n}\right)$ $=\mathbb{P}(\delta_{B}(\mathcal{B}_{0},\hat{\mathcal{B}})\leq 5z_{n,\alpha})$
$=\mathbb{P}\left(\delta_{B}\left(PH_{k}(m_{1}),\widehat{PH}^{\epsilon}_{k}(m_{1})\right)\leq 5z_{n,\alpha}\right)$
$\geq 1-\alpha,$ (By Theorem 3.1 and Equation (4.2))
Given a $(1-\alpha)$-confidence set $\mathcal{C}_{n}$ for persistence barcodes, we can use the significance measure discussed in [23] to declare a bar to be significant. In particular, a bar $a,b]\in\hat{\mathcal{B}}$ is significant at $\alpha$ level of significance, if
$|a-b|>5\sqrt{2}z_{n,\alpha}.$ (4.3)
4.1 Bandwidth Selection
We provide an appropriate choice of the kernel bandwidth $h\equiv h(n)$ to infer structures for a finite sample size $n$. Note that, for any arbitrary bandwidth $h$, the sample size $n$ must be large enough to ensure convergence of the estimation procedure from Theorem 3.1. For example, for $h=0.1$, we require $n>10^{6}$ to ensure convergence from Theorem 3.1 provided that $C_{\epsilon/2}$ is not a sufficiently large quantity. This suggests choosing $h$ appropriately, depending on the sample size $n$ and the unknown population quantities involved in Theorem 3.1.
Given the unknown population quantities $M_{\epsilon}$ and $C_{\epsilon}$, we choose $h=h_{l}$ to infer structures at a $\alpha$ level of significance such that $h_{l}$ solves
$3M_{\epsilon}n\exp\left({-nh^{6}_{l}C_{\epsilon/2}}\right)=\alpha.$ (4.4)
In view of Equation 4.2, note that for all $h\geq h_{l}$, $z_{n,\alpha}=\epsilon$. In other words, by Theorem 3.1, $5\epsilon$ is $(1-\alpha)$-th quantile of the sampling distribution of $\delta_{B}$ $(\widehat{PH}^{\epsilon}_{k}(m_{1}),PH_{k}(m_{1}))$, if the chosen bandwidth $h\geq h_{l}$. On the other hand, for $h\in(0,h_{l})$, $z_{n,\alpha}>\epsilon$ such that $5z_{n,\alpha}$ is $(1-\alpha)$-th quantile of the sampling distribution of $\delta_{B}(\widehat{PH}^{\epsilon}_{k}(m_{1}),PH_{k}(m_{1}))$. This implies that the confidence set $\mathcal{C}_{n}$ is wider for $h\in(0,h_{l})$, and sharper for $h\geq h_{l}$, for a finite sample size $n$.
Now, we discuss the estimation of the unknown population quantities $M_{\epsilon}$ and $C_{\epsilon}$, from the observed data $(Y_{1},X_{1}),\ldots,(Y_{n},X_{n})$. In this context, we discuss only the estimation of unknown quantities $\delta$, $p_{min}$ and $\gamma_{1}$ involved in $C_{\epsilon}$ (see Equation (3.10)) and $M_{\epsilon}$, since the estimation of other unknown quantities is straightforward.
It is straightforward to estimate $\delta$ from the chosen kernel as $\min K(X_{i})$. However, $\min K(X_{i})$ could be a very small quantity close to 0 as $K$ is bounded, which is not desirable as $\delta>0$. In such a situation, we simply take a reasonably larger value than 0, say $\delta=0.1$. Therefore, in what follows, we take $\delta=\max(0.1,\min K(X_{i}))$ to perform data analysis. One can use kernel density estimates of the regressors to estimate $p_{min}$. However, due to kernel artifacts at boundaries, the estimate of $p_{min}$ is close to 0. Note that $p_{min}>0$, and the smaller value will make $C_{\epsilon}$ close to 0, which obstructs the validity of Theorem 3.1. Therefore, we assume that $p_{min}\geq 0.1$ and use this lower bound, which will be sufficient for our purpose.
Next, to estimate $\gamma_{1}$ and $M_{\epsilon}$, we propose the following algorithm. Consider the definition of derivative for $x_{1}<x_{2}\in\mathbb{R}$:
$m_{1}(x_{1})=\lim_{x_{2}\downarrow x_{1}}\frac{m(x_{2})-m(x_{1})}{x_{2}-x_{1}}\approx(m(x^{*})-m(x_{1}))\eta^{-1},$ (4.5)
where $\eta>0$ such that $x_{2}-x_{1}\approx\eta$, and $x^{*}\in(x_{1},x_{1}+\eta)$.
Then, given the noisy observations $Y_{1},\ldots,Y_{n}$ of $m$ at the points $X_{1},\ldots,X_{n}$, respectively. The procedure of estimating derivative values is articulated in the following algorithm.
Algorithm 1 Derivative Estimation Procedure
1:Input: Data $\{(Y_{1},X_{1}),\ldots,(Y_{n},X_{n})\}$.
2:Step 1: Sorting. Sort the regressors in increasing order $X_{(1)},\dots,X_{(n)}$.
3:Step 2: Threshold Selection. Define the threshold: $\eta=Median\{X_{(2)}-X_{(1)},\ldots,X_{(n)}-X_{(n-1)}\}$
4:Step 3: Filtration. Filter $X_{i}$’s and $Y_{i}$’s as follows: $X_{\eta}=\{X_{(i)}:X_{(i+1)}-X_{(i)}<\eta,i=2,\ldots,n\}$, and $Y_{\eta}=\{Y_{(i)}:X_{(i)}\in X_{\eta}\}$.
5:Step 4: First-Order Differences. For points in $X_{\eta}$ and $Y_{\eta}$, compute: $D_{j}=\frac{Y_{(j+1)}-Y_{(j)}}{X_{(j+1)}-X_{(j)}},\quad j=1,\dots,N$, $N=|X_{\eta}|=|Y_{\eta}|$.
6:Step 5: Cleaning. Mitigate noise in $D_{j},j=1,\ldots,N$ by defining: $\hat{Y}^{(1)}=\{D_{j}:D_{j}\in[Q(0.25)-1.5IQR,Q(0.75)+1.5IQR]\}$, where $Q(p)$ denotes $p$-th empirical quantile of $\{D_{j}:j=1,\ldots,N\}$, and IQR=Q(0.75)Q(0.25)IQR=Q(0.75)-Q(0.25).
7:Output: Derivative estimates $\hat{Y}^{(1)}$.
Now, using the above algorithm, the estimates of $\gamma_{1}$ and $M_{\epsilon}$ are defined as:
$\hat{\gamma}_{1}:=\min_{i}\hat{Y}^{(1)}_{i}\text{ and }\hat{M}_{\epsilon}:=\left\lceil\hat{N}_{\epsilon}-\frac{\hat{\gamma}_{1}}{2\epsilon}\right\rceil\text{, where }\hat{N}_{\epsilon}:=\max_{i}\lceil\hat{Y}^{(1)}_{i}/2\epsilon\rceil,\hat{Y}^{(1)}_{i}\in\hat{Y}^{(1)}.$ (4.6)
Note that using derivative estimates as $\frac{Y_{i}-Y_{i-1}}{X_{i}-X_{i-1}}$ is too noisy in general (see, e.g., [17]) that may affect the results adversarially.
Remark 4.1
The main idea of Algorithm 1 is motivated by Equation (4.5), which suggests approximating the true derivative only for those points in the domain for which there exists $x^{*}$ in a small neighborhood. In essence, this suggests avoiding approximating $m_{1}$ in sparse regions of the domain for a good approximation by Equation (4.5). We avoid these sparse regions by defining a suitable threshold $\eta$ in Step 2, where we have taken $\eta=Median\{X_{(2)}-X_{(1)},\ldots,X_{(n)}-X_{(n-1)}\}$. However, $\eta$ can be suitably chosen based on the data such that $\eta>0$ and $X_{\eta}\neq\varnothing$ in Step 3. In general, $\eta=Median\{X_{(2)}-X_{(1)},\ldots,X_{(n)}-X_{(n-1)}\}$ may not be a good choice, as it may be 0 for certain data sets. For instance, in Section 6.1 for cars data $Median\{X_{(2)}-X_{(1)},\ldots,X_{(n)}-X_{(n-1)}\}=0$. Therefore, we define $\eta=mean\{X_{(2)}-X_{(1)},\ldots,X_{(n)}-X_{(n-1)}\}$ for this data set which is >> 0 and works perfectly.
Note that $X_{\eta}$ in Step 3 provides suitable values in the domain from the observed data for which $X_{(i+1)}$ plays the same role for $X_{(i)}$ as $x^{*}$ plays for $x_{1}$ in Equation (4.5), for all $i=2,\ldots,n$. Thus, for a suitably chosen $\eta$, the ratio of first-order differences of $Y\in Y_{\eta}$ (see Step 3) and $X\in X_{\eta}$ (see Step 3) provides good estimated values of $m_{1}$ by Equation (4.5), at the points in $X\in X_{\eta}$. However, due to randomness in $X_{i}$’s, these estimated values can still be noisy. Therefore, we mitigate noise in the estimated derivative values by cleaning them in Step 5.
We implement Algorithm 1 in simulation and real data analysis to estimate $\gamma_{1}$ and $M_{\epsilon}$ in the following sections. It is evident from simulation and real data analysis results in Section 5 and Section 6 that Algorithm 1 works well to infer true structures underlying the data based on Theorem 3.1 and Equation (4.2) or Equation (4.4). Moreover, we provide an explicit implementation of Algorithm 1 in Section 8.5 for the functions used in Section 5 to examine whether it provides reasonably good estimates of the true derivative.